OpenStudy (anonymous):

how does (x+h)(x^2+2xh+h^2)=2x^2+4xh+2h^2

OpenStudy (anonymous):

2= x + h

OpenStudy (anonymous):

Don't understand why but I see how that would work. so why does x+h=2?

OpenStudy (anonymous):

(x+h)(x^2+2xh+h^2)=2x^2+4xh+2h^2 factor out 2 on right side (x+h)(x^2+2xh+h^2)= 2(x^2+2xh+h^2) divide both side by (x^2+2xh+h^2) so (x+h)=2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you're welcome. :)

OpenStudy (anonymous):

So if you have (x+h)(x^2+2xh+h^2) how do you get the answer?